# On the Fourier Coefficients of Critical Gaussian Multiplicative Chaos

### v1

@@ Abstract.
We continue the study of the Fourier coefficients of Gaussian multiplicative chaos (GMC) recently initiated by Garban and Vargas [16]. We show that the Fourier coefficients $c_{n}$ of critical GMC on the unit interval converge to zero in probability as the frequency $n$ tends to infinity, by establishing tightness for $\{(\log n)|c_{n}|\}_{n\geq 1}$.
@@ 1. Introduction
$K(x,y)=\mathbb{E}[X(x)X(y)]=\log_{+}\frac{1}{|x-y|}+f(x,y),\quad\log_{+}(x)=\max(x,0)$
for some smooth function $f:D\times D\longrightarrow\mathbb{R}$. It therefore cannot be defined point-wise and is instead a distribution (in the sense of Schwartz) with probability one. To rigorously define its exponential, one typically considers some regularisation $X_{\epsilon}$ of $X$ to eliminate the logarithmic singularity and defines $\mu_{\gamma}$ as the weak limit of
(1.2) $\mu_{\gamma,\epsilon}(\mathrm{d}x)=e^{\gamma X_{\epsilon}(x)-\frac{\gamma^{2}}{2}\mathbb{E}X_{\epsilon}(x)^{2}}\mathrm{d}x$
as $\epsilon\to 0$, which is known to be non-degenerate if $\gamma<\gamma_{c}=\sqrt{2d}$ [37, 5, 39]. At $\gamma=\gamma_{c}$, a non-trivial limit can still be obtained by considering the so-called derivative martingale [13]
(1.3) $\mu_{\gamma_{c},\epsilon}(\mathrm{d}x)=\big(\hskip-2.00749pt-\hskip-2.00749ptX_{\epsilon}(x)+\sqrt{2}\log(1/\epsilon)\big)e^{\gamma X_{\epsilon}(x)-\frac{\gamma^{2}}{2}\mathbb{E}X_{\epsilon}(x)^{2}}\mathrm{d}x$
or by multiplying (1.2) by a divergent factor of $\sqrt{\log(1/\epsilon)}$ [34, 19]. Both of these procedures yield the same limit up to a multiplicative constant, namely critical GMC, which is non-atomic and supported on a set of Hausdorff measure zero [35]. In this paper, our primary focus will be to study the asymptotic behavior of the Fourier coefficients of this measure.
@@ 1.1. Harmonic analysis of multifractal measures
In a recent work [16], Garban and Vargas initiated the study of the Fourier coefficients of (one-dimensional) GMC, proving in particular that the latter is Rajchman if $\gamma<\sqrt{2}$  [^1] To be precise, they prove this for the GMC arising from the Gaussian free field on the unit circle.. This resolved a longstanding open problem, and together with [15] was the first result of this kind for random multifractal measures. They also supplemented their result by a series of open problems that have since garnered some attention.
For instance, define the Fourier dimension of a measure $\mathrm{dim}_{F}(\mu)$ to be the largest $s$ for which $\hat{\mu}(n)\leq C/n^{s/2}$ for some constant $C>0$, and let $\mu_{\gamma}$ denote GMC on the unit circle with parameter $\gamma$. Using a fourth moment calculation, Garban and Vargas proved that $\mathrm{dim}_{F}(\mu)\geq(1-\gamma^{2}/2)/2$ if $\gamma<1/\sqrt{2}$, and conjectured that
@@
By contrast, the picture for critical GMC is far murkier. It is well known the Fourier dimension of a measure is bounded above by its Hausdorff dimension, which implies that $\mathrm{dim}_{F}(\mu_{\gamma_{c}})=0$ when $\mu_{\gamma_{c}}$ is a critical GMC measure (i.e.  the in-probability limit of (1.3) as $\epsilon\to 0$). To the best of our knowledge, there are no further results in this direction, and the main open problem in [16] was to determine whether or not critical GMC is Rajchman.
@@ 1.2. Main results
In what follows, let $\mu=\mu_{\sqrt{2}}$ be the critical Gaussian multiplicative chaos measure arising from a $\star$–scale invariant field $X$ on $[0,1]$ (defined in Section 2), meaning
$\mu(\mathrm{d}\theta):=\lim_{t\to\infty}\mu_{t}(\mathrm{d}\theta),\quad\text{ where }\mu_{t}(\mathrm{d}\theta):=\sqrt{t}e^{\sqrt{2}X_{t}(\theta)-\mathbb{E}X_{t}(\theta)^{2}}\mathrm{d}\theta,$
where the limit is in probability, in the space of Radon measures on $[0,1]$. Let $c_{n}$ denote the $n$-th Fourier coefficient of $\mu$. Our main result establishes tightness for $\{(\log n)|c_{n}|\}_{n\geq 1}$, which implies that $c_{n}\to 0$ in probability as $n\to\infty$.
@@
For any $\varepsilon>0$, there exists a $K=K(\varepsilon)>0$ such that
$\sup_{n\geq 1}\mathbb{P}\big\{(\log n)|c_{n}|>K\big\}<\varepsilon.$
The $(\log n)^{-1}$ scaling stems from what can loosely be described as a freezing phenomenon, which takes effect as soon as $\gamma\geq 1/\sqrt{2}$ and contributes logarithmic terms in $n$. This is explained in Section 1.3, which also contains a heuristic explanation of why we believe our bound to be sharp—that is, $|c_{n}|$ is indeed of the order of $1/\log n$.
Our proof is based on a second moment estimate for $c_{n}$ on a good event. While the approach is similar in spirit to Berestycki [5] and Lacoin’s [23] constructions of subcritical and critical GMC, respectively, the presence of an oscillating factor in the integral defining $c_{n}$ here introduces significant complications. We also require a more delicate notion of “good” points than in [5, 23], that uses insights from the study of the maximum of log-correlated Gaussian fields, following the works of Bramson–Ding–Zeitouni [7] and Madaule [28]. An overview of the proof is given in Section 1.3.
@@ 1.3. Proof elements
Let $X$ be a $\star$–scale invariant field as above, and consider the multiplicative chaos measure $\mu_{\gamma}:=\lim_{t\to\infty}\mu_{t,\gamma}$ associated to $X$ for $\gamma\in(1/\sqrt{2},\sqrt{2}]$, where
(1.7) $\mu_{t,\gamma}(\mathrm{d}\theta):=\big(\sqrt{t}\cdot\mathbf{1}(\gamma=\sqrt{2})+\mathbf{1}(\gamma\neq\sqrt{2})\big)e^{\gamma X_{t}(\theta)-(\gamma^{2}/2)t}.$
@@
which is finite if and only if $\gamma^{2}<1$ (the so-called $L^{2}$ phase” of GMC). For larger $\gamma$, the expectation is inflated by points $\theta$ where $X_{t}(\theta)$ is atypically large at many scales $t$. To circumvent this, one usually restricts the integral to “good” points $\theta\in\mathcal{G}$ for which $(X_{t}(\theta),t\geq 0)$ is well-behaved [^2] This is the main idea in [5, 23], for instance. At $\gamma=\sqrt{2}$, this is achieved by working on the event that $\{\forall\theta\in[0,1]:\theta\in\mathcal{G}\}$—on which the bound
(1.8) $\iint\mu_{\sqrt{2},t}(\mathrm{d}\theta_{1})\mu_{\sqrt{2},t}(\mathrm{d}\theta_{2})\leq\iint\mathbf{1}(\theta_{1},\theta_{2}\in\mathcal{G})\mu_{\sqrt{2},t}(\mathrm{d}\theta_{1})\mu_{\sqrt{2},t}(\mathrm{d}\theta_{2})$
@@
is that one cannot directly bound as in (1.8), due to the factor $e^{in(\theta_{2}-\theta_{1})}$. In particular, the set $\mathcal{G}$ must be picked so as to guarantee the positivity and real-valuedness of
(1.10) $\iint e^{in(\theta_{2}-\theta_{1})}\mathbf{1}\big(\theta_{1},\theta_{2}\in\mathcal{G}\big)\mu_{\sqrt{2},t}(\mathrm{d}\theta_{1})\mu_{\sqrt{2},t}(\mathrm{d}\theta_{2}).$
@@
$\mathbb{E}\big[e^{\sqrt{2}(X_{t}(\theta_{1})+X_{t}(\theta_{2}))}\mathbf{1}(\theta_{1},\theta_{2}\in\mathcal{G})\big]$
as a function of the gap $\Delta=(\theta_{2}-\theta_{1})$. This further limits the choice of $\mathcal{G}$ to one for which the event $\{\theta_{1},\theta_{2}\in\mathcal{G}\}$ is relatively simple, while being restrictive enough to make the expectation of (1.10) of the correct order (at the very least, compensating for the additional $\sqrt{t}$ normalization present at criticality (1.7)).
We now explain where the $(\log n)^{-1}$ in Theorem 1.1 comes from, and why we believe it to be sharp. To begin with, a variant of the argument above reveals that the dominant contribution to (1.9) comes from $\log|\Delta|^{-1}\asymp\log n$. For concreteness, consider the contribution from $\Delta\in[n^{-100},n^{-1}]$, which is
(1.11) $\asymp te^{-2t}\iint_{|\Delta|\in[n^{-100},n^{-1}]}e^{\sqrt{2}(X_{t}(\theta_{1})+X_{t}(\theta_{2}))}\mathrm{d}\theta_{1}\mathrm{d}\theta_{2}.$
@@
$X_{s}(\theta_{1})\approx X_{s}(\theta_{2}),\quad\forall s\leq r_{\Delta}:=\log|\Delta|^{-1},$
with high probability (see, e.g., Lemma 2.1), and it follows that (1.11) is
$\approx te^{-2t}\iint e^{2\sqrt{2}X_{r_{\Delta}}(\theta_{1})}e^{\sqrt{2}(X_{t}(\theta_{1})-X_{r_{\Delta}}(\theta_{1}))}e^{\sqrt{2}(X_{t}(\theta_{2})-X_{r_{\Delta}}(\theta_{2}))}\mathrm{d}\theta_{1}\mathrm{d}\theta_{2}.$
Furthermore, $e^{\sqrt{2}(X_{t}(\theta_{i})-X_{r_{\Delta}}(\theta_{i}))}$ will typically be of the order of $e^{t-r_{\Delta}}/\sqrt{t-r_{\Delta}}$, which will allow us to bound the integral above by
(1.12) $\int_{n^{-100}}^{n^{-1}}e^{-2r_{\Delta}}\bigg(\int_{0}^{1}e^{2\sqrt{2}X_{r_{\Delta}}(\theta_{1})}\mathrm{d}\theta_{1}\bigg)\mathrm{d}\Delta$
as $t\to\infty$ (noting the change of variables and that $r_{\Delta}$ is bounded in $t$).
For large $n$, we recognize the integral over $\theta_{1}$ as nothing but the (approximate, unnormalized) total mass of a supercritical chaos measure, this time of parameter $\tilde{\gamma}=2\sqrt{2}$. [^3] The emergence of such a chaos was also found by Garban and Vargas (see Theorem 1.3 in [16]), where it is subcritical since they only consider $2\gamma<\sqrt{2}$. Recalling that $r_{\Delta}=\log|\Delta|^{-1}$, this is known to have order
$(\log|\Delta|^{-1})^{-(3\tilde{\gamma}/(2\sqrt{2}))}e^{r_{\Delta}(\tilde{\gamma}^{2}/2-(\tilde{\gamma}/\sqrt{2}-1)^{2})}=e^{3r_{\Delta}}(\log|\Delta|^{-1})^{-3}$
@@ 1.4. Conjectures for subcritical chaos
$\int_{n^{-100}}^{n^{-1}}e^{-\gamma^{2}r_{\Delta}}\bigg(\frac{e^{r_{\Delta}(2\gamma^{2}-(\sqrt{2}\gamma-1)^{2})}}{(\log|\Delta|^{-1})^{3\gamma/\sqrt{2}}}\bigg)\mathrm{d}\Delta.$
@@ 2. Preliminary estimates
(2.1) $\forall\theta_{1}<\theta_{2},\quad\mathbb{E}[X_{t}(\theta_{1})X_{s}(\theta_{2})]=K_{s\land t}(\theta_{2}-\theta_{1})=\int_{1}^{e^{s\land t}}\frac{k(u(\theta_{2}-\theta_{1}))}{u}\mathrm{d}u,$
for some fixed kernel $k\in C^{1}(\mathbb{R})$ associated to $X$, where $s\land t=\min(s,t)$. For convenience, we will assume that $k(0)=1$, $k$ is even and that it vanishes outside of the unit ball around the origin. The following properties then hold by construction:
@@
In particular, we have the following description for the joint law of $X_{t}$ at two different points.
Lemma 2.1 (Lemma 16 in [13]).
For any $\theta_{1}<\theta_{2}$ (both in $[0,1]$) and any $t_{0}\geq 0$, the law of $(X_{s}(\theta_{1}))_{s\leq t_{0}}$ admits the decomposition,
$X_{s}(\theta_{1})=P_{s}^{\theta_{2},\theta_{1}}+Z_{s}^{\theta_{2},\theta_{1}}$
where
$P_{s}^{\theta_{2},\theta_{1}}=-\int_{0}^{s}K_{u}^{\prime\prime}(\theta_{2}-\theta_{1})X_{u}(\theta_{2})\mathrm{d}u+K_{s}^{\prime}(\theta_{2}-\theta_{1})X_{s}(\theta_{2})$
is measurable with respect to $\sigma((X_{s}(\theta_{2})_{s\leq t_{0}})$, and $(Z_{s}^{\theta_{2},\theta_{1}})_{0\leq s\leq t_{0}}$ is a centered Gaussian field independent from the latter, with covariance kernel
$\mathbb{E}\big[Z_{s^{\prime}}^{\theta_{2},\theta_{1}}Z_{s}^{\theta_{2},\theta_{1}}\big]=s\land s^{\prime}-\int_{0}^{s\land s^{\prime}}K_{u}^{\prime}(\theta_{2}-\theta_{1})^{2}\mathrm{d}u.$
This essentially means that for any $\theta_{1},\theta_{2}$, $X_{k}(\theta_{1})$ and $X_{k}(\theta_{2})$ are almost perfectly correlated up to time $k=\log 1/|\theta_{1}-\theta_{2}|$ (and evolve independently after that time, as per property (3) above). For our purposes, we will only need the following estimates reflecting this correlation structure.
@@
Let $k^{\prime}||_{\infty}=\sup_{x\in[-1,1]}|k^{\prime}(x)|$. Fix $t>0$, $r\in[0,t]$ and $s\in[0,t-r]$. Then the following estimates hold uniformly in $\Delta\in[0,e^{-r+1}]$,
[…]
@@
Assume for now that $\Delta\leq e^{-r}$. Then by expanding $k$,
[…]
For the second claim, we make the change of variables $v=e^{-r}u$ to write
$K_{s+r}(\Delta)-K_{r}(\Delta)=\int_{1}^{e^{s}}\frac{k(e^{r}\Delta v)}{v}\mathrm{d}v=\int_{1}^{e^{s}\land(\Delta e^{r})^{-1}}\frac{k(e^{r}\Delta v)}{v}\mathrm{d}v,$
where in the second equality, we used the fact that $k$ is supported on [1,1][-1,1]. Expanding $k$ as in Equation (2.6) yields
[…]
and (2.5) then follows since $\Delta\leq e^{-r}$. The claim for $\Delta\geq e^{-r}$ follows from Equations (2.2) and (2.5) for $\Delta=e^{-r}$, and the estimate
[…]
For any times $t>r>0$, we adopt the notation
(2.7) $X_{t}^{(r)}:=X_{t+r}-X_{r},$
noting that the field $(X_{t}^{(r)}(\theta))_{\theta\in[0,1]}$ is $\star$-scale invariant with kernel $k^{(r)}(x):=k(e^{r}x)$ since
@@
Our argument will require some control on the large values of $X_{r_{n}}$ and $(X_{t}^{(r_{n})})_{t\geq 0}$, where $r_{n}$ is a carefully chosen time depending on the frequency $n$ of the Fourier coefficient being studied (see Proposition 4.3). Following [13, 28], the typical size of $\max_{\theta\in[0,1]}X_{t}(\theta)$ fluctuates around
@@
as made precise by the following proposition.
Proposition 2.3 (Theorem 1.1 in [1]).
Let $m(t)$ be as in Equation (2.8). Then the family
$\Big(\max_{\theta\in[0,1]}X_{t}(\theta)-m(t)\Big)_{t\geq 1},$
is tight.
This is what one expects from any log–correlated field (see, e.g. [4]). In fact, one can typically go further and show that the paths $s\mapsto X_{s}(\theta),s\leq t$ of points $\theta$ for which $X_{t}(\theta)$ is near the maximum cannot fluctuate too wildly. Instead, they display almost linear growth while remaining under a suitably chosen barrier function. This phenomenon was first leveraged by Bramson [9] to study the maximal particle of branching Brownian motion, and underpins most of the works that followed on the extrema of other log–correlated fields [7, 8, 2, 11].
In what follows, we will use the following upper bound essentially due to Lacoin [23].
Proposition 2.4 (Proposition 2.4 in [23]).
Let $(X_{t})_{t\geq 0}$ be defined as in Equation (2.1). Then almost surely,
$\sup_{t\geq e}\sup_{\theta\in[0,1]}\bigg(X_{t}(\theta)-\sqrt{2}t+\frac{\log t}{2\sqrt{2}}-\frac{4\log\log t}{\sqrt{2}}\bigg)<\infty.$
It follows that with high probability, the fields $(X_{t}(\theta))_{t\geq 0}$ remain below
@@
at every point $\theta\in[0,1]$ in the following sense.
Corollary 2.5.
Let $U(t)$ be defined as in Equation (2.9). Then
$\mathbb{P}\Big(\forall\theta\in[0,1],\forall t\geq 0,X_{t}(\theta)\leq U(t)+A\Big)\underset{\text{$A\to\infty$}}{\longrightarrow}1.$

### v2

@@ Abstract.
We continue the study of the Fourier coefficients of Gaussian multiplicative chaos (GMC) recently initiated by Garban and Vargas [15]. We show that if $\{c_{n}\}_{n\geq 1}$ are the Fourier coefficients of critical GMC on the unit interval, then $(\log n)^{\alpha}c_{n}$ converges to zero in probability as $n$ tends to infinity for any $\alpha<1/4$.
@@ 1. Introduction
$K(x,y)=\mathbb{E}[X(x)X(y)]=\log_{+}\frac{1}{|x-y|}+f(x,y),\quad\log_{+}(x)=\max(\log x,0)$
for some smooth function $f:D\times D\longrightarrow\mathbb{R}$. It therefore cannot be defined pointwise and is instead a distribution (in the sense of Schwartz) with probability one. To rigorously define its exponential, one typically regularizes $X$ to obtain a family $(X_{t})_{t}$ of (pointwise defined) Gaussian fields converging to $X$ as $t\to\infty$, and defines $\mu_{\gamma}$ as the weak limit of
(1.2) $\mu_{\gamma,t}(\mathrm{d}x)=e^{\gamma X_{t}(x)-\frac{\gamma^{2}}{2}\mathbb{E}X_{t}(x)^{2}}\mathrm{d}x$
as $t\to\infty$. This limit is known to be non-degenerate if $\gamma<\gamma_{c}=\sqrt{2d}$ [36, 4, 38]. At $\gamma=\gamma_{c}$, a non-trivial limit can still be obtained by considering the so-called derivative martingale [12]
(1.3) $\mu_{\gamma_{c},t}(\mathrm{d}x)=\big(\hskip-2.00749pt-\hskip-2.00749ptX_{t}(x)+\gamma_{c}\mathbb{E}[X_{t}(x)^{2}]\big)e^{\gamma_{c}X_{t}(x)-\frac{\gamma_{c}^{2}}{2}\mathbb{E}X_{t}(x)^{2}}\mathrm{d}x,$
or by multiplying (1.2) by a divergent factor known as the Seneta-Heyde normalization [33, 18], which depends on the regularization procedure. Both of these procedures yield the same limit up to a multiplicative constant, namely critical GMC, which is non-atomic and supported on a set of Hausdorff measure zero [34]. In this paper, our primary focus will be to study the asymptotic behavior of the Fourier coefficients of this measure.
@@ 1.1. Harmonic analysis of multifractal measures
In a recent work [15], Garban and Vargas initiated the study of the Fourier coefficients of (one-dimensional) GMC, proving in particular that this measure is Rajchman if $\gamma<\sqrt{2}$  [^1] To be precise, they prove this for the GMC arising from the Gaussian free field on the unit circle.. This resolved a longstanding open problem, and together with [14] was the first result of this kind for random multifractal measures. They also supplemented their result by a series of open problems that have since garnered some attention.
For instance, define the Fourier dimension of a measure $\mathrm{dim}_{F}(\mu)$ to be the largest $s$ for which $\hat{\mu}(n)\leq Cn^{-s/2}$ for some constant $C>0$, and let $\mu_{\gamma}$ denote GMC on the unit circle with parameter $\gamma$. Using a fourth moment calculation, Garban and Vargas proved that $\mathrm{dim}_{F}(\mu)\geq(1-\gamma^{2}/2)/2$ if $\gamma<1/\sqrt{2}$, and conjectured that
@@
By contrast, the picture for critical GMC is far murkier. It is well-known that the Fourier dimension of a measure is bounded above by its Hausdorff dimension, which implies that $\mathrm{dim}_{F}(\mu_{\gamma_{c}})=0$ when $\mu_{\gamma_{c}}$ is a critical GMC measure (i.e., the in-probability limit of (1.3) as $t\to\infty$). To the best of our knowledge, there are no further results in this direction, and the main open problem in [15] was to determine whether or not critical GMC is Rajchman.
@@ 1.2. Main results
In what follows, let $\mu=\mu_{\sqrt{2}}$ be the critical Gaussian multiplicative chaos measure arising from a $\star$–scale invariant field $X$ on $[0,1]$ (defined in Section 2), meaning that if $(X_{t})_{t\geq 0}$ is a sequence of $\star$-scale cutoff approximations to $X$,
The limit is in probability in the space of Radon measures on $[0,1]$. Let $c_{n}$ denote the $n$-th Fourier coefficient of $\mu$. Our main result is the following.
@@
For any $\alpha<1/4$, the sequence $((\log n)^{\alpha}c_{n})_{n\geq 1}$ converges to zero in probability as $n$ tends to infinity.
The polylogarithmic decay stems from what can loosely be described as a freezing phenomenon, which takes effect as soon as $\gamma\geq 1/\sqrt{2}$. This is explained in Section 1.3, which also contains a heuristic explanation of why we believe the true order of magnitude of $|c_{n}|$ to be closer to $(\log n)^{-1}$.
Our proof is based on a second moment estimate for $c_{n}$ on a good event. While the approach is similar in spirit to Berestycki [4] and Lacoin’s [22] constructions of subcritical and critical GMC, respectively, the presence of an oscillating factor in the integral defining $c_{n}$ here introduces significant complications. In particular, we require a more delicate notion of “good” points than in [4, 22], at which the underlying field $X_{t}(\theta)$ is controlled at scales $t\geq r_{n}$, where $r_{n}$ is a function of the frequency $n$ of the Fourier coefficient at hand. Doing so instead of controlling all scales makes the problem more tractable, at the cost of powers of $(\log n)$. An overview of the proof is given in Section 1.3.
@@ 1.3. Proof elements
Let $X$ be a $\star$–scale invariant field as above, and consider the multiplicative chaos measure $\mu_{\gamma}:=\lim_{t\to\infty}\mu_{\gamma,t}$ associated to $X$ for $\gamma\in(1/\sqrt{2},\sqrt{2}]$, where
(1.7) $\mu_{\gamma,t}(\mathrm{d}\theta):=\big(\sqrt{t}\cdot\mathbf{1}(\gamma=\sqrt{2})+\mathbf{1}(\gamma\neq\sqrt{2})\big)e^{\gamma X_{t}(\theta)-(\gamma^{2}/2)t}.$
@@
which is finite if and only if $\gamma^{2}<1$ (the so-called $L^{2}$ phase” of GMC). For larger $\gamma$, the expectation is inflated by points $\theta$ where $X_{t}(\theta)$ is atypically large at many scales $t$. To circumvent this, one usually restricts the integral to “good” points $\theta\in G$ for which $(X_{t}(\theta),t\geq 0)$ is well-behaved [^2] This is the main idea in [4, 22], for instance.. At $\gamma=\sqrt{2}$, this is achieved by working on the event that $\{\forall\theta\in[0,1]:\theta\in{G}\}$, on which the bound
(1.8) $\iint\mu_{\sqrt{2},t}(\mathrm{d}\theta_{1})\mu_{\sqrt{2},t}(\mathrm{d}\theta_{2})\leq\iint\mathbf{1}(\theta_{1},\theta_{2}\in{G})\mu_{\sqrt{2},t}(\mathrm{d}\theta_{1})\mu_{\sqrt{2},t}(\mathrm{d}\theta_{2})$
@@
is that one cannot directly bound as in (1.8), due to the factor $e^{in(\theta_{2}-\theta_{1})}$. In particular, the set $G$ must be picked so as to guarantee the positivity and real-valuedness of
(1.10) $\iint e^{in(\theta_{2}-\theta_{1})}\mathbf{1}\big(\theta_{1},\theta_{2}\in{G}\big)\mu_{\sqrt{2},t}(\mathrm{d}\theta_{1})\mu_{\sqrt{2},t}(\mathrm{d}\theta_{2}).$
@@
$\mathbb{E}\big[e^{\sqrt{2}(X_{t}(\theta_{1})+X_{t}(\theta_{2}))}\mathbf{1}(\theta_{1},\theta_{2}\in{G})\big]$
as a function of the gap
$\Delta=(\theta_{2}-\theta_{1}).$
This further limits the choice of G{G} to one for which the event $\{\theta_{1},\theta_{2}\in{G}\}$ is relatively simple, while being restrictive enough to make the expectation of (1.10) of the correct order (at the very least, compensating for the additional $\sqrt{t}$ normalization present at criticality (1.7)).
We now explain why we believe that the order of $|c_{n}|$ is typically closer to $(\log n)^{-1}$. To begin with, a variant of the argument above reveals that the dominant contribution to (1.9) comes from $\log|\Delta|^{-1}\asymp\log n$. For concreteness, consider the contribution to $|c_{n,t}|^{2}$ from $\Delta\in[n^{-100},n^{-1}]$, which is
(1.11) $\asymp te^{-2t}\iint_{|\theta_{2}-\theta_{1}|\in[n^{-100},n^{-1}]}e^{\sqrt{2}(X_{t}(\theta_{1})+X_{t}(\theta_{2}))}\mathrm{d}\theta_{1}\mathrm{d}\theta_{2}.$
@@
$X_{s}(\theta_{1})\approx X_{s}(\theta_{2}),\quad\forall s\leq r(\Delta),$
with high probability (see, e.g., Lemma 16 in [12]), where $r(\Delta):=\log|\Delta|^{-1}$ is the branching time. It follows that (1.11) is
$\approx te^{-2t}\iint e^{2\sqrt{2}X_{r(\Delta)}(\theta_{1})}e^{\sqrt{2}(X_{t}(\theta_{1})-X_{r(\Delta)}(\theta_{1}))}e^{\sqrt{2}(X_{t}(\theta_{2})-X_{r(\Delta)}(\theta_{2}))}\mathrm{d}\theta_{1}\mathrm{d}\theta_{2}.$
Furthermore, $e^{\sqrt{2}(X_{t}(\theta_{i})-X_{r(\Delta)}(\theta_{i}))}$ will typically be of the order of $e^{t-r(\Delta)}/\sqrt{t-r(\Delta)}$, which will allow us to bound the integral above by
(1.12) $\int_{n^{-100}}^{n^{-1}}e^{-2r(\Delta)}\bigg(\int_{0}^{1}e^{2\sqrt{2}X_{r(\Delta)}(\theta_{1})}\mathrm{d}\theta_{1}\bigg)\mathrm{d}\Delta$
as $t\to\infty$ (noting the change of variables and that $r(\Delta)$ is bounded in $t$).
For large $n$, we recognize the integral over $\theta_{1}$ as nothing but the (approximate, unnormalized) total mass of a supercritical chaos measure, this time of parameter $\tilde{\gamma}=2\sqrt{2}$. [^3] The emergence of such a chaos was also found by Garban and Vargas (see Theorem 1.3 in [15]), where it is subcritical since they only consider $2\gamma<\sqrt{2}$. Recalling that $r(\Delta)=\log|\Delta|^{-1}$, this is known to have order
$(\log|\Delta|^{-1})^{-(3\tilde{\gamma}/(2\sqrt{2}))}e^{r(\Delta)(\tilde{\gamma}^{2}/2-(\tilde{\gamma}/\sqrt{2}-1)^{2})}=e^{3r(\Delta)}(\log|\Delta|^{-1})^{-3}$
@@ 1.4. Conjectures for subcritical chaos
$\int_{n^{-100}}^{n^{-1}}e^{-\gamma^{2}r(\Delta)}\bigg(\frac{e^{r(\Delta)(2\gamma^{2}-(\sqrt{2}\gamma-1)^{2})}}{(\log|\Delta|^{-1})^{3\gamma/\sqrt{2}}}\bigg)\mathrm{d}\Delta.$
@@ 1.5. Organisation and notation
1.6. Acknowledgements
We would like to thank Zehua He for their careful reading of the manuscript, and in particular for pointing out an issue with a previous version of this work.
@@ 2. Preliminary estimates
(2.1) $\forall\theta_{1}<\theta_{2},\quad\mathbb{E}[X_{t}(\theta_{1})X_{s}(\theta_{2})]=K_{s\land t}(\theta_{2}-\theta_{1}):=\int_{1}^{e^{s\land t}}\frac{k(u(\theta_{2}-\theta_{1}))}{u}\mathrm{d}u,$
for some fixed seed function $k\in C^{2}(\mathbb{R})$ associated to $X$, where $s\land t=\min(s,t)$. For convenience, we will assume that $k(0)=1$, $k$ is even and that it vanishes outside of the unit ball around the origin. The following properties then hold by construction:
@@
An exact description for the joint law of $(X_{t}(\theta_{1}),X_{t}(\theta_{2}))_{t}$ can be found in [12], Lemma 16, which roughly states that $X_{k}(\theta_{1})$ and $X_{k}(\theta_{2})$ are almost perfectly correlated up to time $k=\log 1/|\theta_{1}-\theta_{2}|$ (after which they evolve independently, as per property (3) above). For our purposes, we will only need the following estimate reflecting this correlation structure.
@@
Let $\|k^{\prime}\|_{\infty}=\sup_{x\in[-1,1]}|k^{\prime}(x)|$. Fix $s>0$. Then uniformly in $\Delta\in[-1,1]\setminus\{0\}$,
(2.2) $K_{s}(\Delta)=s\land(\log|\Delta|^{-1})+O(\|k^{\prime}\|_{\infty}).$
@@
Using the fact that $k$ is supported on [1,1][-1,1] and $k(0)=1$,
(2.3) $K_{s}(\Delta)=\int_{1}^{e^{s}\land|\Delta|^{-1}}\frac{k(u|\Delta|)}{u}\mathrm{d}u=\int_{1}^{e^{s}\land|\Delta|^{-1}}\frac{\mathrm{d}u}{u}+\int_{1}^{e^{s}\land|\Delta|^{-1}}\frac{k(u|\Delta|)-k(0)}{u}\mathrm{d}u.$
The first term is equal to the main term in (2.2). By the fundamental theorem of calculus,
$|k(u|\Delta|)-k(0)|\leq\int_{0}^{u|\Delta|}|k^{\prime}(x)|\mathrm{d}x\leq\|k^{\prime}\|_{\infty}u|\Delta|$
for any $u\in[1,e^{s}\land|\Delta|^{-1}]$. The second term in (2.3) is therefore $O(\|k^{\prime}\|_{\infty})$ by the triangle inequality. ∎
For any $t,r>0$, we define the shifted field
(2.4) $\big(X_{t}^{(r)}(\theta)\big)_{\theta\in[0,1]}:=\big((X_{t+r}-X_{r})(\theta)\big)_{\theta\in[0,1]},$
noting that $(X_{t}^{(r)}(\theta))_{\theta\in[0,1]}$ is $\star$-scale invariant with kernel $k^{(r)}(x):=k(e^{r}x)$ since
@@
In particular, $t\mapsto X_{t}^{(r)}(\theta)$ is a standard Brownian motion for any $r>0$ and $\theta\in[0,1]$.
Our argument will require some control on the large values of $(X_{t}(\theta))_{\theta\in[0,1]}$ at multiple scales $t$. For fixed, large $t$, the typical size of $\max_{\theta\in[0,1]}X_{t}(\theta)$ is known to fluctuate around
@@
(see [12, 28]), which is what one expects from any log–correlated field (see, e.g. [3]). In fact, one can typically go further and show that the paths $s\mapsto X_{s}(\theta),s\leq t$, for which $X_{t}(\theta)$ is close to the maximum cannot fluctuate too wildly. Instead, they display approximately linear growth while remaining under a suitably chosen barrier function. This phenomenon was first leveraged by Bramson [6] to study in the maximal displacement of branching Brownian motion, and underpins much of the subsequent work on extrema of log–correlated fields [7, 8, 1, 10].
For our purposes, we will only require the following upper bound of this type, due to Lacoin ([22], Proposition 2.6).
Proposition 2.2.
Let $(X_{t})_{t\geq 0}$ be defined as in Equation (2.1) and
@@
Then as $A\to\infty$,
$\mathbb{P}\bigg(\exists{t\geq 0},\exists{\theta\in[0,1]},X_{t}(\theta)>U(t)+A\bigg)\to 0.$
Remark 2.3.
While [22] assumes that the seed function $k$ is in $C^{\infty}(\mathbb{R})$, the proof of Proposition 2.6 therein only requires that $k\in C^{2}(\mathbb{R})$.